OpenStudy (anonymous):

what is the derivative of x^2x or for that matter x^nx?

OpenStudy (anonymous):

shouldnt this just come on to be 2x(x) -> 2x^2? x^nx -> nx x ?

OpenStudy (phi):

is this \[ x^{nx} \] ?

OpenStudy (anonymous):

correct

OpenStudy (phi):

I think it's a bit tricky y= x^(nx) take the natural log (ln) of both sides ln(y)= ln ( x^(nx)) ln(y)= nx* ln(x) now take the derivative 1/y dy/dx = nx (d/dx ln(x)) + ln(x)* d/dx (nx) replace 1/y with x^(-nx) and complete the derivative on the right side x^(-nx) dy/dx = nx*(1/x) + n*ln(x) dy/dx= n(1+ln(x))x^(nx)

OpenStudy (anonymous):

what if i dont have this set equal to y just the plain derivative of x^nx or is that even possible?

OpenStudy (phi):

the stuff on the right hand side is the "plain derivative" setting it equal to y just makes it easier to type up

OpenStudy (anonymous):

that works then, because what i am asked here is to evaluate a limit with working oh L'Hopital's rule \[\lim_{x \rightarrow 0^-} x^{2x}\]

OpenStudy (anonymous):

actually sorry its \[\lim_{x \rightarrow 0^+}\]

OpenStudy (anonymous):

the answer for this problem is 1 as the limit approaches 0+

OpenStudy (anonymous):

also say for example, i have another question - taking the derivative of \[(\tan x)^{\cos x}\]

OpenStudy (anonymous):

\[(1+\frac{ 1 }{ x })^{\ln x}\]

OpenStudy (phi):

L'Hopital's rule requires \( \frac{\infty}{\infty} \) or \( \frac{0}{0} \) to use it with \(x^{2x} \) I would rewrite it as \[\lim_{x \rightarrow 0+}x^{2x}= e^{\lim_{x \rightarrow 0+} 2x \ln(x)}\] now rewrite x lnx as \[ \frac{\ln x}{x^{-1}} \] This is now in a form where you can apply L'Hopital's rule d/dx ln x = 1/x d/dx x^-1 = -x^2 the ratio is -x and \[ \lim_{x->0+} -x =0\] we get \[ e^{\lim_{x \rightarrow 0+} 2x \ln(x)}= e^0=1\]

OpenStudy (anonymous):

Thank you!